OpenStudy (anonymous):

) Determine the equation for the data, and give the values of the y-intercept and the slope. (Note:In bio-chemistry this is known as a Lineweaver-Burk plot, and the y-intercept is related to the maximum speed of the reaction.

OpenStudy (anonymous):

y-axis x-axis Amount of H2O2 Reaction Speed (amount/second) 1.96 4.75 ×10^−5 1.31 4.03 ×10^−5 0.98 3.51 ×10^−5 0.65 2.52 ×10^−5 0.33 1.44 ×10^−5 0.16 0.585 ×10^−5

OpenStudy (aaronq):

 If you want a Lineweaver-Burk plot, the plot the inverses, \(\dfrac{1}{v}~vs~\dfrac{1}{[H_2O_2]}\). \(y-int=\dfrac{1}{V_{max}}\) \(x-int=-\dfrac{1}{K_m}\)